(a) Explain how the mark-up of price over marginal cost depends on the elasticity of demand. Why can this mark- up be viewed as a measure of monopoly power? P-MC 4. (b) Explain what is meant by the deadweight loss from monopoly power? How might the government limit monopoly power through price regulation? Use diagrams to illustrate your answer.
Answer #1

a) Note that we have mark up = (P - MC)/P. This gives mark up = 1 - MC/P. Now that we have MC = P(1 - 1/e) we see that MC/P is 1 - 1/e. Hence mark up becomes 1 - 1 + 1/e. This indicates that mark up is 1/e where e is the absolute value of elasticity of demand. Thus, mark up percentage is the reciprocal of absolute value of price elasticity of demand. This is a measure of monopoly power because it indicates the divergence between price and marginal cost. If firm has market power, it can charge a price higher than marginal cost. The greater the market power, the larger will be the difference between price and marginal cost implying a larger mark up close to 1.

b) Deadweight loss of a monopoly is the loss of efficiency that results when the monopolist charges a higher price and produces a lower quantity than the competitive market. This is shown as a region between the competitive output and monopoly output bounded by demand curve and marginal cost curve.

Government can reduce the monopoly power by regulating the pricing of monopolist. This is done by forcing the monopoly to charge a price equal to marginal cost or average cost. In both the cases there is a reduction in the monopoly power because now monopoly is producing a higher level of output and is charging a lower price.

Price Figure illustrates the case of a unregulated monopoly and a regulated one. Unregulated monopoly would be using marginal cost and marginal revenue curves to find the profit maximizing output. This would then be used to find the relevant price on the demand function. This is shown by the monopoly price Pm and monopoly quantity Qm. These are the results of an unregulated market. MC Pm Government fixes the price that this monopoly charges according TC to the equation P MC to ensure market efficiency. Firms in a regulated monopoly produce where price and marginal cost are equal. This is shown by Pr as regulated price and Qr as regulated quantity ATC Consumer surplus is now increased from area APmB to APrE Producer surplus is reduced from PmBFIK to PrEFIK Profits are reduced for this regulated monopoly Quantity